jsa:

Aisha wants to make two quilts, each with the same area. The first quilt will be square with sides s feet long. The second quilt will be a rectangle with a width that is half the length of a side of the square quilt and a length that is 6 feet longer than a side length of the square quilt. Which quadratic equation can be used to find s, the side length of the square quilt? s2 = (s + 6) s2 = (s)(s + 6) s2 = (6s) s2 = (s)(6s)

Extrinix:

So we need to know what the equation for a square is, which is \(\sf{w \times l}\), and in this case it’s the same length, so you can just do \(\sf{s^2}\) Now, a rectangle is the same, \(\sf{w \times l}\) but, in this case it’s asking us to use one variable, \(\sf{s}\), and we’re given a value of the length, +6 So given the fact that the rectangle is \(\sf{w \times l+6}\), what do you think the answer is?

Extrinix:

@surjithayer wrote:
\[s^2=\frac{ s }{2 }\times (s+6)\]
How does that help the user?